OpenStudy (anonymous):

can you help me w/ this? indefinite integral of u(1+u)^(2/3) du

OpenStudy (anonymous):

Make a substitution s=u-1 and ds=du. The integral becomes: integral of (s-1)s^(2/3)ds integrate this by distributing and using power rule to get 3/8*s^(8/3)-3/5*s^(5/3)+C back substitute our u back in, and we get: 3/8*(u+1)^(8/3)-3/5*(u+1)^(5/3)+C

OpenStudy (anonymous):

thanks.. It helps a bit... It was actually a part of a larger problem...

OpenStudy (anonymous):

The substitution should be s=u+1. Oops.